solve using the substitution method:
2(x-y)=3+x,
x=3y+4

Answer by user_dude2008(1862)

2(x-y)=3+x

2(3y+4-y)=3+3y+4

2(2y+4)=3y+7

4y+8=3y+7

4y-3y=7-8

y=-1

x=3y+4=3(-1)+4=-3+4=1

Answers: x=1, y=-1